OpenStudy (anonymous):

How can I find (h,k) from the equation g(x) = 2x^2 −16x + 15

OpenStudy (anonymous):

you want to write it like \[g(x)=a(x-h)^2+k\] right?

OpenStudy (anonymous):

Yeah, but how would I get it to that point?

OpenStudy (anonymous):

\[h=-\frac{b}{2a}\] which in this case is \[-\frac{-16}{2\times 2}=4\]

OpenStudy (anonymous):

so it is \[g(x)=2(x-4)^2+k\] the easiest way to find \(k\) is to replace \(x\) by \(4\) i.e. find \(g(4)\)

OpenStudy (anonymous):

in the second form \[g(4)=2(4-4)^2+k=k\] use the first form and compute \[g(4)=2\times 4^2-16\times 4+15\]